OpenStudy (anonymous):

How do I solve -4j - (-5)/-7 = 13?

OpenStudy (anonymous):

Well first off, -4j - (-5) can be changed to -4j +5 because minus negative means add, also the way it's set up, you could multiply the negative to -5 meaning it 5. To remove the fraction. Multiply everything by 7. Can you solve it from here?

OpenStudy (anonymous):

Okay so it's -4j + 5 = -91?

OpenStudy (anonymous):

wolframalpha.com

OpenStudy (anonymous):

Well let me ask first, that is (-4j-(-5))/-7, correct? If so, yes, but that's not solved yet.

OpenStudy (anonymous):

I did -4j +5 - 5 = -91 - 5 which got me -4j = -96 = 24? I think I got it.

OpenStudy (anonymous):

That's correct :)

OpenStudy (anonymous):

Ah, thanks so much! Fractions render me useless :D

OpenStudy (anonymous):

Always remember you can multiply them out, or you could of even done it as (1/7)*(-4j-(-5))=13 if it helps.

OpenStudy (anonymous):

okay, I'll remember that.